$SPAD/src/input intheory.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool intheory.output
)set message test on
)set message auto off
)clear all
--S 1 of 22
div144 := divisors(144)
--R
--R
--R   (1)  [1,2,3,4,6,8,9,12,16,18,24,36,48,72,144]
--R                                                          Type: List(Integer)
--E 1
--S 2 of 22
#(div144)
--R
--R
--R   (2)  15
--R                                                        Type: PositiveInteger
--E 2
--S 3 of 22
reduce(+,div144)
--R
--R
--R   (3)  403
--R                                                        Type: PositiveInteger
--E 3
--S 4 of 22
numberOfDivisors(144)
--R
--R
--R   (4)  15
--R                                                        Type: PositiveInteger
--E 4
--S 5 of 22
sumOfDivisors(144)
--R

--R
--R   (5)  403
--R                                                        Type: PositiveInteger
--E 5
--S 6 of 22
f1(n) == reduce(+,[moebiusMu(d) * numberOfDivisors(quo(n,d)) for d in divisors(n)])
--R
--R                                                                   Type: Void
--E 6
--S 7 of 22
f1(200)
--R
--R   Compiling function f1 with type PositiveInteger -> Integer
--R
--R   (7)  1
--R                                                        Type: PositiveInteger
--E 7
--S 8 of 22
f1(846)
--R
--R
--R   (8)  1
--R                                                        Type: PositiveInteger
--E 8
--S 9 of 22
f2(n) == reduce(+,[moebiusMu(d) * sumOfDivisors(quo(n,d)) for d in divisors(n)])
--R
--R                                                                   Type: Void
--E 9
--S 10 of 22
f2(200)
--R
--R   Compiling function f2 with type PositiveInteger -> Integer
--R
--R   (10)  200
--R                                                        Type: PositiveInteger
--E 10
--S 11 of 22
f2(846)
--R
--R
--R   (11)  846
--R                                                        Type: PositiveInteger
--E 11

--S 12 of 22
fibonacci(25)
--R
--R
--R   (12)  75025
--R                                                        Type: PositiveInteger
--E 12
--S 13 of 22
[fibonacci(n) for n in 1..15]
--R
--R
--R   (13)  [1,1,2,3,5,8,13,21,34,55,89,144,233,377,610]
--R                                                          Type: List(Integer)
--E 13
--S 14 of 22
fib(n) == reduce(+,[binomial(n-1-k,k) for k in 0..quo(n-1,2)])
--R
--R                                                                   Type: Void
--E 14
--S 15 of 22
fib(25)
--R
--R   Compiling function fib with type PositiveInteger -> Integer
--R
--R   (15)  75025
--R                                                        Type: PositiveInteger
--E 15
--S 16 of 22
[fib(n) for n in 1..15]
--R
--R
--R   (16)  [1,1,2,3,5,8,13,21,34,55,89,144,233,377,610]
--R                                                          Type: List(Integer)
--E 16
--S 17 of 22
legendre(3,5)
--R
--R
--R   (17)  - 1
--R                                                                Type: Integer
--E 17
--S 18 of 22
legendre(23,691)

--R
--R
--R   (18)  - 1
--R                                                                Type: Integer
--E 18
--S 19 of 22
h(d) == quo(reduce(+, [jacobi(d,k) for k in 1..quo(-d, 2)]), 2 - jacobi(d,2))
--R
--R                                                                   Type: Void
--E 19
--S 20 of 22
h(-163)
--R
--R   Compiling function h with type Integer -> Integer
--R
--R   (20)  1
--R                                                        Type: PositiveInteger
--E 20
--S 21 of 22
h(-499)
--R
--R
--R   (21)  3
--R                                                        Type: PositiveInteger
--E 21
--S 22 of 22
h(-1832)
--R
--R
--R   (22)  26
--R                                                        Type: PositiveInteger
--E 22
)spool
)lisp (bye)

———-

References^{[1] nothing}

7